How many ways can the letters A, B, C, D, E, and F be arranged in a row 4 at a time if neither A, nor C can be first or third and D cannot be fourth? Repetition is permitted.
Answer by stanbon(75887)

How many ways can the letters A, B, C, D, E, and F be arranged in a row 4 at a time if neither A, nor C can be first or third and D cannot be fourth? Repetition is permitted.
----
_,_,_,_
===============
1st letter: 4 ways
2nd letter: 6 ways
3rd letter: 4 ways
4th letter: 5 ways
--------------------------
Total # of arrangements: 4*6*4*5 = 30*16 = 480 ways
========================
Cheers,
Stan H.
============